A dairy wishes to mix 1,000 pounds of milk containing 8% butterfat. If the mixture is to be made from milk containing 5% butterfat and cream containing 20% butterfat, how many pounds of each is needed?
Found 2 solutions by richwmiller, josgarithmetic:
Answer by richwmiller(17219)


a+b=1000,
0.05*a+0.2*b=0.08*1000
a=1000-b
0.05*(1000-b)+0.2*b=80
50-0.05b+0.2*b=80
0.15*b=30
b=200
a=1000-b
a=800 lb. at 5%
b=200 lb. at 20%
check
0.05*800+0.2*200=0.08*1000
40+40=80
80=80
ok
codemx1

Answer by josgarithmetic(39617)